The meat department in a grocery store needs 200lb of hamburger that is 80% lean beef. How many pounds of 78% lean and 85% lean musted be mixed?
Help!! I wish I was better at word problems. Thanking you in advance for your help.

Answer by nerdybill(7384)

The meat department in a grocery store needs 200lb of hamburger that is 80% lean beef. How many pounds of 78% lean and 85% lean musted be mixed?
.
Let x = amount (pounds) of 78% lean
then
200-x = amount (pounds) of 85% lean
.
.78x + .85(200-x) = .80(200)
.78x + 170-.85x = 160
170-.07x = 160
-.07x = -10
x = 142.86 pounds (78% lean)
.
85% lean:
200-x = 200-142.86 = 57.14 pounds